# LIA: Privacy-Preserving Data Quality Evaluation in Federated Learning Using a Lazy Influence Approximation

### v3

@@ Abstract
In Federated Learning, it is crucial to handle low-quality, corrupted, or malicious data. However, traditional data valuation methods are not suitable due to privacy concerns. To address this, we propose a simple yet effective approach that utilizes a new influence approximation called "lazy influence" to filter and score data while preserving privacy. To do this, each participant uses their own data to estimate the influence of another participant’s batch and sends a differentially private obfuscated score to the central coordinator. Our method has been shown to successfully filter out biased and corrupted data in various simulated and real-world settings, achieving a recall rate of over $>90\%$ (sometimes up to $100\%$) while maintaining strong differential privacy guarantees with $\varepsilon\leq 1$.
1 Introduction
@@
Imagine a collaborative project involving multiple hospitals across the globe, each with its own repository of chest X-ray scans of patients. The goal is to use FL to train a robust ML model capable of diagnosing heart conditions from these X-rays [51]. However, each hospital adheres to different data collection protocols. E.g., Hospital A may use a different scanner making it incompatible with the rest of the data. If included, it can lead to a large drop in accuracy [42]. More insidiously, Hospital B may habitually include textual annotations directly on the scans for internal communication. The model may rely on these shortcuts instead of looking at the actual scan [48].
@@
A clean way of quantifying the quality of data point(s) is via the notion of influence [35, 11]. Intuitively, influence quantifies the marginal contribution of a data point (or batch of points) on a model’s accuracy. One can compute this by comparing the difference in the model’s empirical risk when trained with and without the point in question. While the influence metric can be highly informative, it is impractical to compute: re-training a model is time-consuming, costly, and often impossible, as participants do not have access to the entire dataset. We propose a simple and practical approximation of the sign of the exact influence (Lazy Influence Approximation) which is based on an estimate of the direction of the model after a small number of local training epochs with the new data.
@@
The second challenge is to limit the information leaked by this approximate influence measure to preserve the privacy of the data. Many approaches to Federated Learning (e.g., [46, 55]) remedy this by combining FL with Differential Privacy (DP) [17, 18, 19, 20, 31, 15, 9], a data anonymization technique that many researchers view as the gold standard [54]. We show how the sign of influence can be approximated in an FL setting while maintaining strong differential privacy guarantees. Specifically, there are two sets of participants’ data that we need to protect: the training and the validation data (see also Section 1.2). We use DP-SGD for this training to ensure datapoint-level local DP [1, 66] To ensure the privacy of the validation data and the influence approximation itself, we employ a differentially private defense mechanism based on the idea of randomized response [59] (inspired by [22]). Together the two mechanisms ensure strong, worst-case differential privacy guarantees while allowing for accurate data filtering. See Figure 1 for an overview of our approach.
1.1 Our Contributions
@@
1.2 High-Level Description of Our Setting
A center $C$ coordinates a set of participants that contribute their data to train a single model (Figure 1). $C$ has a small set of ‘warm-up’ data, which are used to train an initial model $M_{0}$ that captures the desired input/output relation. The model is then updated with the contributions of the participants. We assume that each data holder participant has a set of training points that will be used to improve the model and a set of validation points that will be used to evaluate other participants’ contributions. It must be kept private to prohibit participants from tailoring their contributions to the validation data. Each data holder participant can assume two roles: the role of the contributor ($A$) and the role of the validator ($B$). As a contributor, a participant performs a small number of local epochs to $M_{t}$ – enough to get an estimate of the gradient [^2] The number of local epochs is a hyperparameter. We do not need to train the model fully (e.g., 2-5 epochs). – using a batch of his training data $z_{A,t}$. Subsequently, $A$ sends the updated partial model $M_{t,A}$, with specifically crafted noise to ensure DP, to the validators. The applied noise protects the updated gradient while still retaining information on the usefulness of data. Each validator $B$ uses its validation dataset to approximate the empirical risk of $A$’s training batch (i.e., the approximate influence). This is done by evaluating each validation point and comparing the loss. In an FL setting, we can not re-train the model to compute the exact influence; instead, $B$ performs only a small number of training epochs, enough to estimate the direction of the model (Lazy Influence Approximation). As such, we look at the sign of the approximate influence (and not the magnitude). Each validator aggregates the signs of the influence for each validation point, applies controlled noise to ensure DP, and sends this information to the center. Finally, the center accepts $A$’s training batch if most of the $B$s report positive influence and rejects otherwise (majority voting).
2 Related Work and Discussion
@@
Influence functions are a standard method from robust statistics [11] (see also Section 3), which were recently used as a method of explaining the predictions of black-box models [35]. They have also been used in the context of fast cross-validation in kernel methods and model robustness [43, 10]. While a powerful tool, computing the influence involves too much computation and communication, and it requires access to the training and validation data (see [35] and Section 3). There has also been recent work trying to combine Federated Learning with influence functions [62], though to the best of our knowledge, we are the first to provide a privacy-preserving alternative.
@@
Differential Privacy (DP) [17, 18, 19, 20] has emerged as the de facto standard for protecting the privacy of individuals. Informally, DP captures the increased risk to an individual’s privacy incurred by participating in the learning process. Consider a participant being surveyed on a sensitive topic as a simplified, intuitive example. To achieve differential privacy, one needs a source of randomness; thus, the participant decides to flip a coin. Depending on the result (heads or tails), the participant can reply truthfully or randomly. An attacker can not know if the decision was taken based on the participant’s preference or due to the coin toss. Of course, to get meaningful results, we need to bias the coin toward the actual data. In this simple example, the logarithm of the ratio $Pr[\text{heads}]/Pr[\text{tails}]$ represents the privacy cost (also referred to as the privacy budget), denoted traditionally by $\varepsilon$. Yet, one must be careful in designing a DP mechanism, as it is often hard to practically achieve a meaningful privacy guarantee (i.e., avoid adding a lot of noise and maintain high accuracy) [55, 13]. A variation of DP, instrumental in our context, given the decentralized nature of federated learning, is Local Differential Privacy (LDP) [34, 21]. LDP is a generalization of DP that assumes all communication between the data providers and the Center is public. Hence, we do not rely on the Center to take appropriate privacy measures and instead try limit information leakage in all communication. This allows us to work with untrusted central coordinators. An orthogonal concept is datapoint-level vs. user-level DP [58, 4, 44]. In user-level privacy, we attempt to protect datasets at the level of the data provider, rather than at the level of data-point. While this is typical in the cross-device FL where each data-provider represents the data of a single user, this does not make sense for multi-institutional collaboration [41]. In the setting we consider, each hospital (data provider) aggregates data from many patients. Thus, each user we want to protect is associated with a single data point, making datapoint-level DP more relevant. Thus, we consider datapoint-level local DP. We assume that the participants and the Center are honest but curious, i.e., they don’t actively attempt to corrupt the protocol but will try to learn about each other’s data.
3 Methodology
@@
3.1 Shortcomings of the Exact and Approximate Influence in a FL Setting
@@
3.2 Lazy Influence (LIA): A Practical Influence Metric for Filtering Data in FL Applications
@@
(i) At federated learning round $t$ (model $M_{t}(\theta_{t})$), the contributor participant $A$ performs a small number $k$ of local epochs to $M_{t}$ using a batch of his training data $Z_{A,t}$, resulting in $\tilde{\theta}_{t}^{A}$. $k$ is a hyperparameter. $\tilde{\theta}_{t}^{A}$ is the partially trained model of participant $A$, where most of the layers, except the last one, have been frozen. The model should not be fully trained for two key reasons: efficiency and avoiding over-fitting (e.g., in our simulations, we only performed 1-20 epochs). Furthermore, $A$ adds noise to $\tilde{\theta}_{t}^{A}$ (see Section 3.2.2) to ensure strong, worst-case local differential privacy. Finally, $A$ sends only the last layer (to reduce communication cost) of $\tilde{\theta}_{t}^{A}$ to every other participant.
(ii) Each validator $B$ uses his validation dataset $Z_{val}^{B}$ to estimate the sign of the influence using Equation 1. Next, the validator applies noise to $\mathcal{I}_{LIA}(Z_{val}^{B})$, as described in Section 3.2.3, to ensure strong, worst-case differential privacy guarantees (i.e., keep his validation dataset private).
@@ 12            $B_{i}$: Sends $v^{\prime}$ to $C$ (private majority voting)
3.2.1 Advantages of the proposed lazy influence
@@
3.2.2 Sharing the Partially Updated Model: Privacy, Communication Cost, and Scalability
Each contributor $A$ shares a partially trained model $\tilde{\theta}_{t}^{A}$ (see step (i) of Section 3.2). It is important to stress that A only sends the last layer of the model. This has two significant benefits. First, it minimizes the impact of the differential privacy noise. We follow  [1, 66] to ensure strong datapoint-level local differential privacy guarantees by (i) imposing a bound on the gradient (using a clipping threshold $\Delta$), and (ii) adding carefully crafted Gaussian noise (parameterized by $\sigma$). Second, it reduces the communication overhead (e.g., in our CIFAR simulations, we only send $0.009\%$ of the model’s weights). Moreover, as explained in the Introduction, this communication cost will be incurred only once, when we use our approach as a ‘right of passage’ every time a participant joins the federation. Importantly, in practice one can randomly select a (sufficiently large [^3] With enough votes, the aggregated LIA signs will match the sign of the true influence value in expectation. The size of the validators set is problem specific. See Sections 4, 4.3, and Figure 4.) subset of validators, keeping the communication cost constant as we scale.
@@ 3.2.2 Sharing the Partially Updated Model: Privacy, Communication Cost, and Scalability
3.2.3 Differentially Private Reporting of the Influence
@@
Each validator $B_{i}$ is assigned a random ID number every time they act in that capacity. Moreover, each can split their test data into multiple disjoint sets, and then use a each set only once. [^4] E.g., in our applications, each of these sets required at most 50 points for the vote to be accurate. We obfuscate their influence reports using RAPPOR [22], which results in an $\varepsilon$-differential privacy guarantee [20]. [^5] Note that RAPPOR provides significantly stronger privacy guarantees ($\delta=0$) that what is normally deployed in modern systems. Thus, in practice, the privacy loss $\varepsilon$ will be even lower if we allow $\delta>0$. The obfuscation (permanent randomized response [59]) takes as input the participant’s true influence value $v$ (binary) and privacy parameter $p$, and creates an obfuscated (noisy) reporting value $v^{\prime}$, according to Eq. 2. $p$ is a user-tunable parameter that allows the participants themselves to choose their desired level of privacy, while maintaining reliable filtering. The worst-case privacy guarantee can be computed by each participant a priori, using Eq. 3 [22]. $v^{\prime}=\begin{cases}+1,&\text{with probability $\frac{1}{2}p$}\\ -1,&\text{with probability $\frac{1}{2}p$}\\ v,&\text{with probability $1-p$}\end{cases}$ (2) $\varepsilon=2\ln\left(\frac{1-\frac{1}{2}p}{\frac{1}{2}p}\right)$ (3)
It is important to note that in a Federated Learning application, the center $C$ aggregates the influence sign from a large number of participants. This means that even under really strict privacy guarantees, the aggregated influence signs (which is exactly what we use for filtering) will match the sign of the true value in expectation. This results in high-quality filtering, as we will demonstrate in Section 4.
@@ 4 Evaluation Results
We evaluated the proposed approach on four well-established datasets: CIFAR10 [36], CIFAR100 [36], real-data on Human Activity Recognition (HAR) on edge devices [3, 2], and real-data on diagnosing heart conditions from X-rays [51]. Between these datasets, we evaluate on two different modalities (image data and sensor data), and corruption of both the input X, and the labels y. Due to lack of space, we focus on CIFAR in this section. Similar results were achieved for the other datasets. Please see the Appendix for detailed results (and also the bar plot in Figure 1).
@@
Our evaluation involves a single round of Federated Learning. A small portion of every dataset (around 1%) is selected as the ‘warm-up’ data used by the center $C$ to train the initial model $M_{0}$. Each participant has two datasets: a training batch ($Z_{A}$, see Section 3.2, step (i)), which the participant uses to update the model when acting as the contributor participant, and a validation dataset ($Z_{val}^{B}$, see Section 3.2, step (ii)), which the participant uses to estimate the sign of the influence when acting as a validator participant. As a concrete example, for the CIFAR simulations the ratio of these datasets is $2:1$. The training batch size is 100 (i.e., the training dataset includes 100 points, and the validation dataset consists of 50 points). This means that, e.g., for a simulation with 100 participants, each training batch is evaluated on $50\times(100-1)$ validation points, and that for each training batch (contributor participant $A$), the center collected ($100-1$) estimates of the influence sign (Equation 1). See the Appendix for details on the HAR dataset. We report results when having 30% corrupt participants. Additional results (0%, 10%, 20%, 40%) can be found in the Appendix. For each corrupted batch, we corrupted 100% of the data points (similar results were achieved with 90%). Each simulation was run 8 times. We report average values and standard deviations. The proposed approach is model-agnostic and can be used with any gradient-descent-based ML method.
@@
The main hurdle for FL is that not all data is IID. Heterogeneous data distributions are all but uncommon in the real world. To simulate non-IID data, we used the Dirichlet distribution to split the training dataset as in related literature [28, 40, 27, 68]. This distribution is parameterized by $\alpha$, which controls the concentration of different classes (see the Appendix for a visualization). We report results for $\alpha\rightarrow 0.1$, as in related literature (e.g., [68]), see the Appendix for the rest. This translates to a highly non-IID distribution, e.g., with just 3 classes per participant for the HAR dataset.

### v4

@@ Abstract
In Federated Learning, it is crucial to handle low-quality, corrupted, or malicious data. However, traditional data valuation methods are not suitable due to privacy concerns. To address this, we propose a simple yet effective approach that utilizes a new influence approximation called ”lazy influence” to filter and score data while preserving privacy. To do this, each participant uses their own data to estimate the influence of another participant’s batch and sends a differentially private obfuscated score to the central coordinator. Our method has been shown to successfully filter out biased and corrupted data in various simulated and real-world settings, achieving a recall rate of over $>90\%$ (sometimes up to $100\%$) while maintaining strong differential privacy guarantees with $\varepsilon\leq 1$.
Index Terms:
Federated Learning, Data Filtering, Data Poisoning, Robustness, Privacy.
I Introduction
@@
Imagine a collaborative project involving multiple hospitals across the globe, each with its own repository of chest X-ray scans of patients. The goal is to use FL to train a robust ML model capable of diagnosing heart conditions from these X-rays [51]. However, each hospital adheres to different data collection protocols. E.g., Hospital A may use a different scanner making it incompatible with the rest of the data. If included, it can lead to a large drop in accuracy [42]. More insidiously, Hospital B may habitually include textual annotations directly on the scans for internal communication. The model may rely on these shortcuts instead of looking at the actual scan [48]. At the same time, it is important not to discard data about rare diseases, even if it might appear very different from most other data.
@@
To prevent such a scheme from filtering out valuable data about rare conditions, quality assessment cannot be based on simple data similarity, which would consider them outliers. A clean way of quantifying the quality of data point(s) is via the notion of influence [35, 11]. Intuitively, influence quantifies the marginal contribution of a data point (or batch of points) on a model’s accuracy. One can compute this by comparing the difference in the model’s empirical risk when trained with and without the point in question. While the influence metric can be highly informative, it is impractical to compute: re-training a model is time-consuming, costly, and often impossible, as participants do not have access to the entire dataset. We propose a simple and practical approximation of the sign of the exact influence (Lazy Influence Approximation) which is based on an estimate of the direction of the model after a small number of local training epochs with the new data.
@@
The second challenge is to limit the information leaked by this approximate influence measure to preserve the privacy of the data. Many approaches to Federated Learning (e.g., [46, 55]) remedy this by combining FL with Differential Privacy (DP) [17, 19, 20, 31, 15, 9], a data anonymization technique that many researchers view as the gold standard [54]. We show how the sign of influence can be approximated in an FL setting while maintaining strong differential privacy guarantees. Specifically, there are two sets of participants’ data that we need to protect: the training and the validation data (see also Section I-B). We use DP-SGD for this training to ensure datapoint-level local DP [1, 66] To ensure the privacy of the validation data and the influence approximation itself, we employ a differentially private defense mechanism based on the idea of randomized response [59] (inspired by [22]). Together the two mechanisms ensure strong, worst-case differential privacy guarantees while allowing for accurate data filtering. See Figure 1 for an overview of our approach.
I-A Our Contributions
@@
I-B High-Level Description of Our Setting
A center $C$ coordinates a set of participants that contribute their data to train a single model (Figure 1). $C$ has a small set of ‘warm-up’ data, which are used to train an initial model $M_{0}$ that captures the desired input/output relation. The model is then updated with the contributions of the participants. We assume that each data holder participant has a set of training points that will be used to improve the model and a set of validation points that will be used to evaluate other participants’ contributions. It must be kept private to prohibit participants from tailoring their contributions to the validation data. Each data holder participant can assume two roles: the role of the contributor ($A$) and the role of the validator ($B$). As a contributor, a participant performs a small number of local epochs to $M_{t}$ – enough to get an estimate of the gradient [^2] The number of local epochs is a hyperparameter. We do not need to train the model fully (e.g., 2-5 epochs). – using a batch of his training data $z_{A,t}$. Subsequently, $A$ sends the updated partial model $M_{t,A}$, with specifically crafted noise to ensure DP, to the validators. The applied noise protects the updated gradient while still retaining information on the usefulness of data. Each validator $B$ uses its validation dataset to approximate the empirical risk of $A$’s training batch (i.e., the approximate influence). This is done by evaluating each validation point and comparing the loss. In an FL setting, we can not re-train the model to compute the exact influence; instead, $B$ performs only a small number of training epochs, enough to estimate the direction of the model (Lazy Influence Approximation). As such, we look at the sign of the approximate influence (and not the magnitude). Each validator aggregates the signs of the influence for each validation point, applies controlled noise to ensure DP, and sends this information to the center. Finally, the center accepts $A$’s training batch if most of the $B$s report positive influence and rejects otherwise (majority voting). To recognize the value of data about rare conditions, it is important to use a broad sample of validation data. Following the federated learning idea, we distribute evaluation among many participants. Anyone capable of using the model can be a validator, so there could be many more than contributors.
II Related Work and Discussion
@@
Influence functions are a standard method from robust statistics [11] (see also Section III), which were recently used as a method of explaining the predictions of black-box models [35]. They have also been used in the context of fast cross-validation in kernel methods and model robustness [43, 10]. While a powerful tool, computing the influence involves too much computation and communication, and it requires access to the training and validation data (see [35] and Section III). There has also been recent work trying to combine Federated Learning with influence functions [62], though to the best of our knowledge, we are the first to provide a privacy-preserving alternative.
@@
Differential Privacy (DP) [17, 19, 20] has emerged as the de facto standard for protecting the privacy of individuals. Informally, DP captures the increased risk to an individual’s privacy incurred by participating in the learning process. Consider a participant being surveyed on a sensitive topic as a simplified, intuitive example. To achieve differential privacy, one needs a source of randomness; thus, the participant decides to flip a coin. Depending on the result (heads or tails), the participant can reply truthfully or randomly. An attacker can not know if the decision was taken based on the participant’s preference or due to the coin toss. Of course, to get meaningful results, we need to bias the coin toward the actual data. In this simple example, the logarithm of the ratio $Pr[\text{heads}]/Pr[\text{tails}]$ represents the privacy cost (also referred to as the privacy budget), denoted traditionally by $\varepsilon$. Yet, one must be careful in designing a DP mechanism, as it is often hard to practically achieve a meaningful privacy guarantee (i.e., avoid adding a lot of noise and maintain high accuracy) [55, 13]. A variation of DP, instrumental in our context, given the decentralized nature of federated learning, is Local Differential Privacy (LDP) [34, 21]. LDP is a generalization of DP that assumes all communication between the data providers and the Center is public. Hence, we do not rely on the Center to take appropriate privacy measures and instead try limit information leakage in all communication. This allows us to work with untrusted central coordinators. An orthogonal concept is datapoint-level vs. user-level DP [58, 4, 44]. In user-level privacy, we attempt to protect datasets at the level of the data provider, rather than at the level of data-point. While this is typical in the cross-device FL where each data-provider represents the data of a single user, this does not make sense for multi-institutional collaboration [41]. In the setting we consider, each hospital (data provider) aggregates data from many patients. Thus, each user we want to protect is associated with a single data point, making datapoint-level DP more relevant. Thus, we consider datapoint-level local DP. We assume that the participants and the Center are honest but curious, i.e., they don’t actively attempt to corrupt the protocol but will try to learn about each other’s data.
III Methodology
@@
III-A Shortcomings of the Exact and Approximate Influence in a FL Setting
@@
III-B Lazy Influence (LIA): A Practical Influence Metric for Filtering Data in FL Applications
@@
(i) At federated learning round $t$ (model $M_{t}(\theta_{t})$), the contributor participant $A$ performs a small number $k$ of local epochs to $M_{t}$ using a batch of his training data $Z_{A,t}$, resulting in $\tilde{\theta}_{t}^{A}$. $k$ is a hyperparameter. $\tilde{\theta}_{t}^{A}$ is the partially trained model of participant $A$, where most of the layers, except the last one, have been frozen. The model should not be fully trained for two key reasons: efficiency and avoiding over-fitting (e.g., in our simulations, we only performed 1-20 epochs). Furthermore, $A$ adds noise to $\tilde{\theta}_{t}^{A}$ (see Section III-B2) to ensure strong, worst-case local differential privacy. Finally, $A$ sends only the last layer (to reduce communication cost) of $\tilde{\theta}_{t}^{A}$ to every other participant.
(ii) Each validator $B$ uses his validation dataset $Z_{val}^{B}$ to estimate the sign of the influence using Equation 1. Next, the validator applies noise to $\mathcal{I}_{LIA}(Z_{val}^{B})$, as described in Section III-B3, to ensure strong, worst-case differential privacy guarantees (i.e., keep his validation dataset private).
@@ 12            $B_{i}$: Sends $v^{\prime}$ to $C$ (private majority voting)
III-B1 Advantages of the proposed lazy influence
@@
III-B2 Sharing the Partially Updated Model: Privacy, Communication Cost, and Scalability
Each contributor $A$ shares a partially trained model $\tilde{\theta}_{t}^{A}$ (see step (i) of Section III-B). It is important to stress that A only sends the last layer of the model. This has two significant benefits. First, it minimizes the impact of the differential privacy noise. We follow  [1, 66] to ensure strong datapoint-level local differential privacy guarantees by (i) imposing a bound on the gradient (using a clipping threshold $\Delta$), and (ii) adding carefully crafted Gaussian noise (parameterized by $\sigma$). Second, it reduces the communication overhead (e.g., in our CIFAR simulations, we only send $0.009\%$ of the model’s weights). Moreover, as explained in the Introduction, this communication cost will be incurred only once, when we use our approach as a ‘right of passage’ every time a participant joins the federation. Importantly, in practice one can randomly select a (sufficiently large [^3] With enough votes, the aggregated LIA signs will match the sign of the true influence value in expectation. The size of the validators set is problem specific. See Sections IV, IV-C, and Figure 4.) subset of validators, keeping the communication cost constant as we scale.
@@ III-B2 Sharing the Partially Updated Model: Privacy, Communication Cost, and Scalability
III-B3 Differentially Private Reporting of the Influence
@@
Each validator $B_{i}$ is assigned a random ID number every time they act in that capacity. Moreover, each can split their test data into multiple disjoint sets, and then use a each set only once. [^4] E.g., in our applications, each of these sets required at most 50 points for the vote to be accurate. We obfuscate their influence reports using RAPPOR [22], which results in an $\varepsilon$-differential privacy guarantee [20]. [^5] Note that RAPPOR provides significantly stronger privacy guarantees ($\delta=0$) that what is normally deployed in modern systems. Thus, in practice, the privacy loss $\varepsilon$ will be even lower if we allow $\delta>0$. The obfuscation (permanent randomized response [59]) takes as input the participant’s true influence value $v$ (binary) and privacy parameter $p$, and creates an obfuscated (noisy) reporting value $v^{\prime}$, according to Eq. 2. $p$ is a user-tunable parameter that allows the participants themselves to choose their desired level of privacy, while maintaining reliable filtering. The worst-case privacy guarantee can be computed by each participant a priori, using Eq. 3 [22].
$v^{\prime}=\begin{cases}+1,&\text{with probability $\frac{1}{2}p$}\\ -1,&\text{with probability $\frac{1}{2}p$}\\ v,&\text{with probability $1-p$}\end{cases}$ (2)
$\varepsilon=2\ln\left(\frac{1-\frac{1}{2}p}{\frac{1}{2}p}\right)$ (3)
It is important to note that in a Federated Learning application, the center $C$ aggregates the influence sign from a large number of participants. This means that even under really strict privacy guarantees, the aggregated influence signs (which is exactly what we use for filtering) will match the sign of the true value in expectation. This results in high-quality filtering, as we will demonstrate in Section IV.
@@ IV Evaluation Results
We evaluated the proposed approach on four well-established datasets: CIFAR10 [36], CIFAR100 [36], real-data on Human Activity Recognition (HAR) on edge devices [3, 2], and real-data on diagnosing heart conditions from X-rays [51]. Between these datasets, we evaluate on two different modalities (image data and sensor data), and corruption of both the input X, and the labels y. Due to lack of space, we focus on CIFAR in this section. Similar results were achieved for the other datasets. Please see the supplement (XI) for detailed results.
@@
Our evaluation involves a single round of Federated Learning. A small portion of every dataset (around 1%) is selected as the ‘warm-up’ data used by the center $C$ to train the initial model $M_{0}$. Each participant has two datasets: a training batch ($Z_{A}$, see Section III-B, step (i)), which the participant uses to update the model when acting as the contributor participant, and a validation dataset ($Z_{val}^{B}$, see Section III-B, step (ii)), which the participant uses to estimate the sign of the influence when acting as a validator participant. As a concrete example, for the CIFAR simulations the ratio of these datasets is $2:1$. The training batch size is 100 (i.e., the training dataset includes 100 points, and the validation dataset consists of 50 points). This means that, e.g., for a simulation with 100 participants, each training batch is evaluated on $50\times(100-1)$ validation points, and that for each training batch (contributor participant $A$), the center collected ($100-1$) estimates of the influence sign (Equation 1). See the supplement (XI) for details on the HAR dataset. We report results when having 30% corrupt participants. Additional results (0%, 10%, 20%, 40%) can be found in the supplement XI. For each corrupted batch, we corrupted 100% of the data points (similar results were achieved with 90%). Each simulation was run 8 times. We report average values and standard deviations. The proposed approach is model-agnostic and can be used with any gradient-descent-based ML method.
@@
The main hurdle for FL is that not all data is IID. Heterogeneous data distributions are all but uncommon in the real world. To simulate non-IID data, we used the Dirichlet distribution to split the training dataset as in related literature [28, 40, 27, 68]. This distribution is parameterized by $\alpha$, which controls the concentration of different classes (see the supplement VII-A for a visualization). We report results for $\alpha\rightarrow 0.1$, as in related literature (e.g., [68]), see the supplement XI for the rest. This translates to a highly non-IID distribution, e.g., with just 3 classes per participant for the HAR dataset.